OpenStudy (anonymous):

Find all solutions of the equation in the interval [0,2pi) . sinΘ =5/7 Do not round any intermediate computations, and round your answer(s) to the nearest hundredth.

OpenStudy (anonymous):

arc sin (5/7)= 45.58o and 180- 45.58 = 134.41o or convert to radians.

OpenStudy (anonymous):

how would I convert to radians? 45.580 * pi/180 ?

OpenStudy (anonymous):

nevrermind thanks! :)

OpenStudy (anonymous):

it says its wrong..

OpenStudy (anonymous):

Wrong angle? Wrong procedure? arc sin(5/7) = 45.58... certainly, and similarly at 180o minus this. Puzzling.